OpenStudy (landon34):

if n-3/4=2/3, then which of the following is true?

OpenStudy (landon34):

answers:3n – 3 = 8 3n – 9 = 6 3n – 9 = 8 3n – 3 = 6

OpenStudy (landon34):

the first part of the equation is a fraction.

OpenStudy (landon34):

Hello?